OpenStudy (anonymous):

I think it is C, am i right? Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(0.2)x f(x) = 4(0.98)x f(x) = 0.7(5)x f(x) = 5(4)−x

jimthompson5910 (jim_thompson5910):

why C?

OpenStudy (anonymous):

i really dont know, i think it is also B as well

jimthompson5910 (jim_thompson5910):

Rule: If you have \[\Large f(x) = a*(b)^x\] f(x) will have exponential growth if `b > 1` or f(x) will have exponential decay if `0 < b < 1`

OpenStudy (anonymous):

so it is neither C or B..

jimthompson5910 (jim_thompson5910):

Flip things around If you have \[\Large f(x) = a*(b)^{-x}\] f(x) will have exponential decay if `b > 1` or f(x) will have exponential growth if `0 < b < 1`

jimthompson5910 (jim_thompson5910):

which function has `b > 1` ? and has `x` (not `-x`) in the exponent?

OpenStudy (anonymous):

it has to be A then

jimthompson5910 (jim_thompson5910):

in choice A, what are the values of a and b?

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

for `f(x) = 0.4(0.2)^x`, the value of a and b are a = 0.4 b = 0.2 agreed?

OpenStudy (anonymous):

oh yes, im sorry i miss read what you said.

jimthompson5910 (jim_thompson5910):

so b = 0.2 is b > 1 true?

OpenStudy (anonymous):

i think it is

jimthompson5910 (jim_thompson5910):

0.2 > 1 is true?

jimthompson5910 (jim_thompson5910):

here is a graph to interactively play with https://www.desmos.com/calculator/fhg9iipwt0

jimthompson5910 (jim_thompson5910):

adjust the sliders for 'a' and 'b'. Make sure 'a' stays as a positive number

OpenStudy (anonymous):

a stayed as a positive number @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah make a = 2 or something then adjust the slider for b

OpenStudy (anonymous):

they stayed positive @jim_thompson5910